# Removing the need for ground truth UWB data collection: self-supervised ranging error correction using deep reinforcement learning

### v1

@@ I Introduction
Precise indoor positioning technology has attracted significant research interest in recent years due to its role in overcoming the limitations of global positioning system (GPS) in indoor environments for Internet of Things (IoT) applications such as assistive healthcare systems [1], sports tracking [2], smart logistics [3] and various location-based services [4]. Following this trend, Ultra-Wideband (UWB) technology has seen a surge in interest and become one of the more promising technologies for indoor positioning systems (IPS). UWB IPS can achieve centimeter-level positioning accuracy due to the wide bandwidth ($>$500 MHz) and very short time duration of the pulse (around 2 ns) [5]. While these signal characteristics make UWB more resilient to multipath effects (compared to traditional narrowband techniques such as SigFox, LoRa, Narrowband Internet of Things (NB-IoT), etc. [6, 7]). However, a major remaining challenge is correcting ranging errors caused by this multipath behavior in non-line-of-sight (NLOS) conditions [8], [9]. Current methods to detect and reduce errors caused by NLOS conditions rely mostly on machine/deep learning models trained using large datasets of UWB ranges and raw physical data like the channel impulse response (CIR) [10, 11, 12] or calculated features [7] (e.g. amplitude of the signal, energy, power ratio, etc.) labeled with the true positions. While these approaches can lead to high performance, it comes with two major disadvantages. First, collecting such labeled data requires a tedious labeling effort and dataset collection, which requires specialized equipment and personnel with expertise in UWB positioning and ground truth data collection. Second, the usability is limited by the generalization problem, the accuracy of trained solutions drops severely in unseen environments. The unseen environments have different anchor topologies, different sizes, and different UWB hardware, or contain different types of objects that degrade the performance due to (1) variations in the CIR and (2) different UWB physical layer (PHY) properties. The generalization problem worsens the data collection problem as each unique environment requires the labeling effort to be repeated and even so, the environment may have changed by then. These two disadvantages have previously been addressed using (1) semi-supervised learning [11] [13], to reduce the data collection and (2) transfer learning to enable better performance in unseen environments while using only a few labeled samples [14, 15]. However, all these approaches still require some labeled samples, thus a tedious data collection effort. One other research proposes a self-supervised ranging error correction [16] that does not require ground truth or label collection. It uses classical location approaches to jointly estimate the location and range with a deep network in a Time Difference Of Arrival (TDoA) system. However, the learning here is limited as they do not use signal features to aid and improve the learning process, and they cannot correct separate ranges.
@@ II-A Supervised learning
The authors of [7] propose a feature-based approach with both support vector machine (SVM) regression and a Gaussian process (GP) to form an estimate of the ranging error. The authors of [12] propose an approach using latent variables that encapsulate information from the CIR about both distance and environmental features to then employ variational inference techniques with neural networks to perform approximate inference in a supervised manner. The authors of [10] propose a supervised deep learning approach for UWB ranging error correction. It leverages a probabilistic deep learning architecture by combining variational inference with probabilistic neural networks. The approach uses a variational autoencoder to learn features from the CIR. [11] uses a similar autoencoder approach for feature extraction from the CIR, but the models are trained in a dual-loss fashion to jointly optimize unsupervised autoencoding and supervised prediction. While both [11] and [10] leverage unsupervised pre-training of the autoencoder layers, the key ranging error prediction task is formulated as a supervised learning problem. Here, labeled data is used to train a model to directly map inputs to known target outputs. These papers show that supervised machine learning approaches using both raw physical data (CIR) or features can be used to significantly improve the UWB ranging performance. However, none of them address the problem of data collection or the generalization problem, meaning that real-world usability is limited.
@@ II-B Semi-supervised learning
Paper Self-supervised ML approach
@@
Environment type Input for learning
Final outcome
@@
TWR LOS/NLOS Features, range Ranging error
@@
TWR LOS/NLOS CIR Ranging error
@@
TWR LOS/NLOS CIR Ranging error
@@
[14] Transfer learning TWR LOS/NLOS Features, CIR
@@
Our work RL TWR LOS/NLOS CIR Ranging Error
@@ II-D Self-supervised learning
To the best of our knowledge, [16] is the only self-supervised approach for UWB error mitigation. The authors propose a deep location and ranging correction (DLRC) network to jointly estimate the tag position corrections and distance corrections in a UWB localization system using TDoA. The approach tries to extract the high-level spatial features from ten consecutive ranging measurements received at all anchors in the system, by using the topological information of the UWB system, location loss, and ranging loss are minimized together, allowing training without the ground truth. This approach varies significantly from ours as it was developed for a TDoA system, utilizing measurements from various anchors, without leveraging the valuable information in the CIR data for error correction. This highlights a research gap where our proposed self-supervised reinforcement learning method specifically targets single-range measurements within a TWR ranging system utilizing CIR data and removes the need for ground truth data collection.
@@ III Dataset description
The dataset is collected in an industrial lab environment, which is part of the Industrial Internet of Things (IIoT) testbed [17] of the IDLab research group at Ghent University. The lab is a 240 m2 warehouse environment, representative of many Industry 4.0 use cases. The IIoT testbed consists of an open space area and an area with metal racks, leading to line-of-sight (LOS) and NLOS situations, pictured in Figure 2(a). The environment is equipped with 18 Qualisys Miqus M3 Motion Capture (MOCAP) cameras, capable of tracking hundreds of passive infrared reflective MOCAP markers with a quantified uncertainty in the millimeter range at speeds up to 340 Hz, enabling accurate ground truth determination for evaluation purposes (not used for training in this research). In addition, the MOCAP system is used in combination with a mobile robotic platform to drive repeatable trajectories through the lab. A total of 23 anchors are distributed over the environment, the placement is illustrated in Figure 2(b) with the light blue crosses. The dataset was collected using Wi-PoS devices [18] that carry the Qorvo DW1000 UWB transceivers. During measurement, the CIR information used for learning was captured at the anchor nodes. To capture the data a mobile robot drives around the lab at 0.1 m/s, the trajectory of the robot is shown in Figure 2(b). This trajectory leads to 3463 UWB ranging samples with the different anchors. The ranging method used in the system is called Asymmetric double-sided TWR (ADS-TWR) [19]
@@ III-A Data pre-processing
Second, the RSSI-array is trimmed to 150 samples, 50 samples before and 100 after the estimated first path by the DW1000 using the leading edge algorithm. Lastly, the remaining array is normalized using min-max normalization. This means that the highest value in the array becomes 1 and the lowest 0:
@@ IV-A UWB Localization System
Where t represents the timestamp for each value within the CIR (one CIR has 1016 complex values, corresponding to ±$\pm 10^{-}9s$ each); S is the number of multipath components; $\alpha_{s}$ is the amplitude of the s-th multipath component; $\tau_{s}$ the time delay of the s-th multipath component; $\delta$ the Dirac delta function and n represents the additive white Gaussian noise (AWGN) present in the channel. In NLOS conditions, $fp$ can be severely attenuated and the calculated ToF becomes inaccurate:
@@
With $fp^{\prime}$ the first detected path above the noise floor $n(fp-1)$ and the real first path $fp$ not detected or not yet received. $\tau_{fp^{\prime}-fp}$ is the time difference between the detected first path and the real first path. The ranging error this causes can be calculated as:
@@ V Proposed methodology
$\gamma$ Discount factor, determining weight of target critic
@@
$p_{KF}$ Kalman Filter position
@@
In our methodology, we assume occasional movements of people or vehicles in the environment, which follow sufficiently predictable trajectories. Combining this predictability with filtering, smoothing, and error correction, improvements in error correction are rewarded over time. This is achieved using a RL process that continually enhances the filtered and corrected positions, leading to continuously improving data available for ranging correction.
@@ V-C Data processing for self-supervised reward
As discussed before, the actor determines ranging correction $\widehat{e_{t}}$ and the target actor generates correction $\dot{e_{t}}$ that leads to the corrected target range $\dot{\Delta}_{t}=\widehat{\Delta_{t}}-\dot{e_{t}}$ and this corrected range estimate is used to iteratively improve the range correction and self-generate better labels. $\dot{\Delta}_{t}$ is converted to a position using a Kalman Filter [22].
$p_{KF,t}=\text{Kalman\textunderscore Filter}(\dot{\Delta}_{t})$ (16)
Then added to a circular buffer $C$ of length N (assumed odd) used for smoothing. If the buffer is full when a new $p_{KF,t}$ is added to the buffer together with its associated $\Delta^{\prime}_{t}$, $CIR_{t}$ and $\widehat{e_{t}}$, the oldest value in the buffer is removed and the average position of all positions in the buffer is determined and linked to the value at the middle position of the buffer:
@@
This $p_{avg,m}$ is related to the remaining data at position $m$ in the circular buffer: $\Delta^{\prime}_{m}$, $CIR_{m}$ and $\widehat{e_{m}}$. Finally, to get an improved range estimate, $p_{avg,m}$ is converted back to a range:$\Delta_{avg,m}$, by calculating the Euclidean distance with the anchor $a_{n}$. This value is our current best estimate of the range and is used in the reward function:
@@
The goal of the reward function is to provide a quantitative measure of the success of an agent’s actions in the environment. By shaping the reward function appropriately, we can guide the agent to exhibit the desired behavior, namely improved range accuracy. This reward function gives higher rewards the closer the corrected range of the RL agent is to the current best estimate of the range. Updating the neural network at every time step with one sample would be very inefficient. Therefore, the network is updated on batches of data that are sampled from a replay memory containing experiences $(CIR_{m},\widehat{e_{m}},R_{m})$ generated during the execution of the algorithm. There are several methods to sample from this memory, for this problem we opted for random sampling instead of prioritized sampling as we do not want to overfit on certain experiences or have a lack of diversity in the sampled experiences. An illustration of the complete proposed methodology is shown in Figure 4 and the pseudocode is given in Algorithm 1.
@@
The network architecture of the actor and critic is given in Table III. The actor network ends with scaling the output of a dense layer with a Tanh activation function. The Tanh output is between -1 and 1, which leads to the final output being scaled to -1000 and 1000 which is equal to the action space. The critic network starts similar to the actor network, but the information of the CIR is encapsulated in 4 latent features. These 4 latent features are concatenated with the action selected by the actor. This is then further processed to a final layer with a Tanh activation, meaning that the Q-value is between -1 and 1.
@@
Input (State) (1,150,1) Input (State) (1,150,1)
Conv2D(128,16x1) ReLU (128,150,1) Conv2D(128,16x1) ReLU (128,150,1)
Maxpool(2x1) (128,75,1) Maxpool(2x1) (128,75,1)
Conv2D(64,8x1) ReLU (64, 75, 1) Conv2D(64,8x1) ReLU (64, 75, 1)
Conv2D(32,2x1) ReLU (32,75,1) Conv2D(32,2x1) ReLU (32,75,1)
BatchNorm (32,75,1) BatchNorm (32,75,1)
Dropout 25% (32,75,1) Dropout 25% (32,75,1)
@@ VI-A Baselines and metrics
For performance evaluation, we will use two evaluation metrics: (1) the mean absolute error (MAE) as it encapsulates the performance in a single value and (2) box plots to provide a clear and concise way to see the spread (variability) of ranging errors and thus a more general overview of the performance while also highlighting central tendencies. To evaluate our proposed method, we compare our results against two baselines: the first baseline is the uncorrected UWB performance, and the second is a state-of-the-art supervised CNN-based method [11] trained on the fully labeled dataset. The results of the supervised CNN are not directly adopted from the paper itself, but the developed model has been retrained on the dataset of this research. The range error results of NLOS samples will be shown and discussed separately because of the reduced signal clarity in NLOS situations. For NLOS, the signal propagation between transmitter and receiver is more complex due to the attenuated first path signal power, leading to a more complicated relationship between CIR and error correction. The NLOS situations are the most vital for error correction, as they are prone to the largest ranging errors. Showing the performance in NLOS situations separately provides insight into how the baselines and our proposed RL algorithm perform in the most challenging conditions.

### v2

@@ I Introduction
Precise indoor positioning technology has attracted significant research interest in recent years due to its role in overcoming the limitations of global positioning system (GPS) in indoor environments for Internet of Things (IoT) applications such as assistive healthcare systems [1], sports tracking [2], smart logistics [3] and various location-based services [4]. Following this trend, Ultra-Wideband (UWB) technology has seen a surge in interest and become one of the more promising technologies for indoor positioning systems (IPS). UWB IPS can achieve centimeter-level positioning accuracy due to the wide bandwidth ($>$500 MHz) and very short time duration of the pulse (around 2 ns) [5]. While these signal characteristics make UWB more resilient to multipath effects (compared to traditional narrowband techniques such as SigFox, LoRa, Narrowband Internet of Things (NB-IoT), etc. [6, 7]). However, a major remaining challenge is correcting ranging errors caused by this multipath behavior in non-line-of-sight (NLOS) conditions [8], [9]. Current methods to detect and reduce errors caused by NLOS conditions rely mostly on machine/deep learning models trained using large datasets of UWB ranges and raw physical data like the channel impulse response (CIR) [10, 11, 12] or calculated features [7] (e.g. amplitude of the signal, energy, power ratio, etc.) labeled with the true positions. While these approaches can lead to high performance, it comes with two major disadvantages. First, collecting such labeled data requires a tedious labeling effort and dataset collection, which requires specialized equipment and personnel with expertise in UWB positioning and ground truth data collection. Second, the usability is limited by the generalization problem, the accuracy of trained solutions drops severely in unseen environments. The unseen environments have different anchor topologies, different sizes, and different UWB hardware, or contain different types of objects that degrade the performance due to (1) variations in the CIR and (2) different UWB physical layer (PHY) properties. The generalization problem worsens the data collection problem as each unique environment requires the labeling effort to be repeated and even so, the environment may have changed by then. These two disadvantages have previously been addressed using (1) semi-supervised learning [11] [13], to reduce the data collection and (2) transfer learning to enable better performance in unseen environments while using only a few labeled samples [14, 15]. However, all these approaches still require some labeled samples, thus a tedious data collection effort. One other research proposes a self-supervised ranging error correction [16] that does not require ground truth or label collection. It uses classical location approaches to estimate the location and range with a deep network jointly. However, the learning here is limited as they do not use signal features to aid and improve the learning process, and they cannot correct separate ranges.
@@ II-A Supervised learning
The authors of [7] propose a feature-based approach with both support vector machine (SVM) regression and a Gaussian process (GP) to form an estimate of the ranging error. The authors of [12] propose an approach using latent variables that encapsulate information from the CIR about both distance and environmental features to then employ variational inference techniques with neural networks to perform approximate inference in a supervised manner. The authors of [10] propose a supervised deep learning approach for UWB-ranging error correction. It leverages a probabilistic deep learning architecture by combining variational inference with probabilistic neural networks. The approach uses a variational autoencoder to learn features from the CIR. [11] uses a similar autoencoder approach for feature extraction from the CIR. Still, the models are trained in a dual-loss fashion to optimize unsupervised autoencoding and supervised prediction jointly. While both [11] and [10] leverage unsupervised pre-training of the autoencoder layers, the key ranging error prediction task is formulated as a supervised learning problem. Here, labeled data is used to train a model to directly map inputs to known target outputs. These papers show that supervised machine learning approaches using both raw physical data (CIR) or features can be used to significantly improve the UWB ranging performance. However, none of them address the problem of data collection or the generalization problem, meaning that real-world usability is limited.
@@ II-B Semi-supervised learning
Paper
Self-
supervised
ML approach
@@
Environment type #Anchors Input
Output
@@
TWR LOS/NLOS 5 Features, range Ranging error
@@ 10 CIR
TWR LOS/NLOS 1 CIR Ranging error
@@
TWR LOS/NLOS 19 CIR Ranging error
@@
[14] Transfer learning TWR LOS/NLOS 21 Features, CIR
@@
Our work RL TWR LOS/NLOS 23 CIR Ranging Error
@@ II-D Self-supervised learning
To the best of our knowledge, [16] is the only self-supervised approach for UWB error mitigation. While both our method and [16] aim to improve ranging accuracy, there are significant differences in our approaches, driven by the distinct environments we target. The authors of [16] designed their method for simple, line-of-sight (LOS) environments with few anchors, while our approach addresses complex, large-scale industrial environments with both LOS and NLOS conditions. These environmental differences lead to contrasting algorithm designs. The authors of [16] propose a deep location and ranging correction (DLRC) network that jointly estimates tag position corrections and distance corrections using Time of Arrival (ToA). Their method requires 10 ranges for each anchor for a single error correction. This thus assumes the constant availability of all anchors with a high update rate, otherwise there is no correction available. In contrast, our method uses only one CIR and range per correction, allowing variable anchor availability and update rates typical in large-scale realistic environments. For instance, in our industrial testbed with 23 anchors, [16]’s approach would necessitate determining 230 ranges, leading to significant delays. Our method ensures faster processing by gathering only one range and its corresponding CIR. The information used for learning also differs significantly as the authors of [16] assume minimal tag movement between ranges, limiting the diversity of environmental effects captured. Our use of CIR as input encapsulates detailed information about the transmission channel, allowing our method to correlate the channel state with correction factors more effectively. While both approaches use similar principles in their loss/reward functions, there are notable differences. The method of [16] uses a dual loss function to simultaneously minimize position and range corrections. Our reinforcement learning-based method maximizes cumulative rewards based on the agreement between corrected ranges and Kalman filter positioning. This approach offers continuous adaptation capabilities, allowing our algorithm to adjust to environmental changes in real time – a crucial advantage in dynamic industrial environments.
@@ III Dataset description
The dataset is collected in an industrial lab environment, which is part of the Industrial Internet of Things (IIoT) testbed [17] of the IDLab research group at Ghent University. The lab is a 240 m2 warehouse environment, representative of many Industry 4.0 use cases. The IIoT testbed consists of an open space area and an area with metal racks, leading to LOS and NLOS situations, pictured in Figure 2(a). The environment is equipped with 18 Qualisys Miqus M3 Motion Capture (MOCAP) cameras, capable of tracking hundreds of passive infrared reflective MOCAP markers with a quantified uncertainty in the millimeter range at speeds up to 340 Hz, enabling accurate ground truth determination for evaluation purposes (not used for training in this research). In addition, the MOCAP system is used in combination with a mobile robotic platform to drive repeatable trajectories through the lab. A total of 23 anchors are distributed over the environment, the placement is illustrated in Figure 2(b) with the light blue crosses. The dataset was collected using Wi-PoS devices [18] that carry the Qorvo DW1000 UWB transceivers. During measurement, the CIR information used for learning was captured at the anchor nodes. To capture the data a mobile robot drives around the lab at 0.1 m/s, the trajectory of the robot is shown in Figure 2(b). This trajectory leads to 3463 UWB ranging samples with the different anchors. The ranging method used in the system is called Asymmetric double-sided TWR (ADS-TWR) [19] The same environment can change over time. To evaluate the performance of our proposed algorithm when there are changes in the environment, a second dataset was collected in the same warehouse 6 months later. At that time, there were more goods in the racks, additional clutter in the warehouse (obstacles, boxes, …) and the anchor nodes experienced many small disturbances over time. These combined effects lead to a more challenging environment. This second dataset is smaller and contains 1434 samples.
@@ III-A Data pre-processing
Second, the RSSI-array is trimmed to 150 samples, 50 samples before and 100 after the estimated first path by the DW1000 using the leading edge algorithm. Lastly, the remaining array is normalized using min-max normalization. We subtract the minimum value from each element to make the lowest value zero, and divide by the difference between the maximum and minimum values to scale the highest value to one:
@@ IV-A UWB Localization System
Where t represents the timestamp for each value within the CIR (one CIR has 1016 complex values, corresponding to ±$\pm 10^{-}9s$ each); S is the number of multipath components; $\alpha_{s}$ is the amplitude of the s-th multipath component; $\tau_{s}$ the time delay of the s-th multipath component; $\delta$ the Dirac delta function and $n(t)$ represents the additive white Gaussian noise (AWGN) present in the channel. In NLOS conditions, $fp$ can be severely attenuated and the calculated ToF becomes inaccurate:
@@
With $fp^{\prime}$ the first detected path above the noise floor and earlier multipath or the true first path are not detected. $\tau_{fp^{\prime}-fp}$ is the time difference between the detected first path and the real first path. The ranging error this causes can be calculated as:
@@ V Proposed methodology
$\gamma$ Discount factor, determining the weight of target critic
@@
$p_{EKF}$ Extended Kalman Filter position
@@
In our methodology, we assume occasional movements of people or vehicles in the environment, which follow sufficiently predictable trajectories. Combining this predictability with filtering, smoothing, and error correction, improvements in error correction are rewarded over time. This is achieved using an RL process that continually enhances the filtered and corrected positions, leading to continuously improving data available for ranging correction.
@@ V-C Data processing for self-supervised reward
As discussed before, the actor determines ranging correction $\widehat{e_{t}}$ and the target actor generates correction $\dot{e_{t}}$ that leads to the corrected target range $\dot{\Delta}_{t}=\widehat{\Delta_{t}}-\dot{e_{t}}$ and this corrected range estimate is used to iteratively improve the range correction and self-generate better labels. $\dot{\Delta}_{t}$ is converted to a position using an Extended Kalman Filter (EKF)[22].
$p_{EKF,t}=\text{EKF}(\dot{\Delta}_{t})$ (16)
Then added to a circular buffer $C$ of length N (assumed odd) used for smoothing. The size of the buffer impacts the smoothening and eventually the corrections, it is a trade-off, larger buffer size increases accuracy on straight paths but decreases accuracy in corners and other less predictable movements. If the buffer is full when a new $p_{KF,t}$ is added to the buffer together with its associated $\Delta^{\prime}_{t}$, $CIR_{t}$ and $\widehat{e_{t}}$, the oldest value in the buffer is removed and the average position of all positions in the buffer is determined and linked to the value at the middle position of the buffer:
@@
This $p_{avg,m}$ is related to the remaining data at position $m$ in the circular buffer: $\Delta^{\prime}_{m}$, $CIR_{m}$ and $\widehat{e_{m}}$. The reason for selecting the middle position in the buffer is to ensure that the averaged position, $p_{avg,m}$, is associated with the most representative data points in the buffer. By taking the middle position, you ensure that it is related to the same number of positions before and after. If you associate this average with a data point at the beginning or end of the set, it may not accurately represent the data points around that specific position. This is because the average is biased towards the data points on the side where there are more points. Finally, to get an improved range estimate, $p_{avg,m}$ is converted back to a range:$\Delta_{avg,m}$, by calculating the Euclidean distance with the anchor $a_{n}$. This value is our current best estimate of the range and is used in the reward function:
@@
The goal of the reward function is to provide a quantitative measure of the success of an agent’s actions in the environment. By shaping the reward function appropriately, we can guide the agent to exhibit the desired behavior, namely improved range accuracy. This reward function gives higher rewards the closer the corrected range of the RL agent is to the current best estimate of the range. Updating the neural network at every time step with one sample would be very inefficient. Therefore, the network is updated on batches of data that are sampled from a replay memory containing experiences $(CIR_{m},\widehat{e_{m}},R_{m})$ generated during the execution of the algorithm. There are several methods to sample from this memory, for this problem we opted for random sampling instead of prioritized sampling as we do not want to overfit certain experiences or have a lack of diversity in the sampled experiences. An illustration of the complete proposed methodology is shown in Figure 4 and the pseudocode is given in Algorithm 1.
@@
The network architecture of the actor and critic is given in Table III. The actor network is based on the CNN from [14], a state-of-the-art supervised model. Three convolutional layers are trained to extract local time series features from each CIR. The kernel of size 16 captures more large-scale patterns while the two layers with smaller kernels capture more fine-grained patterns in the data. The progression from 128 to 32 channels is designed to gradually reduce the dimensionality while preserving features. The only difference with [14] is that the actor network ends with scaling the output of a dense layer with a Tanh activation function instead of no activation in the final layer. The Tanh output is between -1 and 1, which leads to the final output being scaled to -1000 and 1000 which is equal to the action space. Not changing the activation function for the actor network leads to training instability and does not result in a good-performing system. Changing the activation function in the supervised CNN also improves the supervised performance. The influence of this change on the supervised CNN is demonstrated in the evaluation.
The critic network starts similar to the actor network, but the information of the CIR is encapsulated in 4 latent features. These 4 latent features are concatenated with the action selected by the actor. This is then further processed to a final layer with a Tanh activation, meaning that the Q-value is between -1 and 1.
@@
Input (State) (1,150) Input (State) (1,150)
Conv1D(128,16) ReLU (128,150) Conv1D(128,16) ReLU (128,150)
Maxpool(2) (128,75) Maxpool(2) (128,75)
Conv1D(64,8) ReLU (64, 75) Conv1D(64,8) ReLU (64, 75)
Conv1D(32,2) ReLU (32,75) Conv2D(32,2) ReLU (32,75)
BatchNorm (32,75) BatchNorm (32,75)
Dropout 25% (32,75) Dropout 25% (32,75)
@@ VI-A Baselines and metrics
For performance evaluation, we will use two evaluation metrics: (1) the mean absolute error (MAE) as it encapsulates the performance in a single value and (2) box plots to provide a clear and concise way to see the spread (variability) of ranging errors and thus a more general overview of the performance while also highlighting central tendencies. To evaluate our proposed method, we compare our results against two baselines: the first baseline is the uncorrected UWB performance, and the second is the state-of-the-art supervised CNN method [14] trained on the fully labeled dataset. The results of the supervised CNN are not directly adopted from the paper itself, but the developed model has been retrained on the dataset of this research. The range error results of NLOS samples will be shown and discussed separately because of the reduced signal clarity in NLOS situations. For NLOS, the signal propagation between transmitter and receiver is more complex due to the attenuated first path signal power, leading to a more complicated relationship between CIR and error correction. The NLOS situations are the most vital for error correction, as they are prone to the largest ranging errors. Showing the performance in NLOS situations separately provides insight into how the baselines and our proposed RL algorithm perform in the most challenging conditions.